For $4.50 a grocer buys a case of fruit which contains 14 dozen. She knows that 4 dozen will spoil before she sells them. At what price per dozen must she sell the good ones to gain 1/3 of the whole cost?
is there any quick way to do this question.
please respond.
Answer #1
(1/3) * $4.50 = $1.50. = Income.

P = $1.50 / (14-4) = $0.15 / doz.
answered by: tod
Answer #2
Actually I believe the answer to this question focuses on a gain, being 1/3 over the initial price of $4.50 (since 4 dozen will spoil before they can be sold, the loss has to be made up). So the income would be: $4.50 * 1.333 = $6.00 And then the price would be: $6.00/(14-4) = $0.60/dozen.
source: Electrical Apprenticeship Exam Prep Guide
answered by: Naomi